OpenStudy (jiteshmeghwal9):

A particle starts from a point with a velocity of 6m,s and moves with an acceleration of -2m/s^2. Show that after 6s the particle will be at the starting point.

OpenStudy (jiteshmeghwal9):

@UnkleRhaukus plz help :)

OpenStudy (unklerhaukus):

\[x(t)=x_0+{v_x}_0\cdot t+\tfrac12a_x\cdot t^2\]

OpenStudy (jiteshmeghwal9):

is this the third equation of motion ?

OpenStudy (jiteshmeghwal9):

@UnkleRhaukus

OpenStudy (unklerhaukus):

its an equation of motion, under constant acceleration

OpenStudy (jiteshmeghwal9):

ok! how this equation can prove the given question ?

OpenStudy (unklerhaukus):

if the object starts at the origin ie x_0 = 0[m], and the initial velocity is v_x_0 = 6[m/s] and the acceleration is a_x = -2[m/s^2]

OpenStudy (unklerhaukus):

what do you get when you substitute t= 6[s] ?

OpenStudy (jiteshmeghwal9):

\[x(t)=x_0+v_{x0}.t+\frac{1}{2}a_x.t^2\]\[x(6)=0+6m/s+\frac{1}{2}-2(6)^2\]

OpenStudy (jiteshmeghwal9):

\[x(6)=6-36\]\[6x=-30\]\[x=-5\]

OpenStudy (unklerhaukus):

you missed t on the middle term

OpenStudy (unklerhaukus):

\[x(6)=0+6m/s\times\color{brown}{6s}+\frac{1}{2}(-2m/s^2)(6s)^2\]

OpenStudy (jiteshmeghwal9):

oh! yes sorry :P

OpenStudy (jiteshmeghwal9):

\[6x=0\]\[x=0\]

OpenStudy (unklerhaukus):

do you mean x(6) the position x, as a a function of time; where the time is 6 seconds

OpenStudy (jiteshmeghwal9):

so it is x(6)=0 only ?

OpenStudy (unklerhaukus):

we could have chosen a different \(x_0\) [ie \(x(0)\)], but the function would still return the same place it started after six seconds.

OpenStudy (jiteshmeghwal9):

ok!

OpenStudy (jiteshmeghwal9):

so now it is the proof right ??